OpenStudy (anonymous):

What is the solution to the inequality √x <9

OpenStudy (anonymous):

this is just wrong

OpenStudy (anonymous):

if \[\sqrt{x}<9\] then \[0<x<81\]

OpenStudy (anonymous):

really? for example \[25>3\] but \[\sqrt{25}=5<9\] yes? and \[-3<-1<3\] but \[\sqrt{-1}\] is not a real number. no, i am right.